# Link Prediction with Untrained Message Passing Layers

### v1

@@ 2. UTSAGE: $h_{v}^{(l)}=\frac{1}{\tilde{d}_{v}}\sum_{u\in\tilde{\mathcal{N}}(v)}h_{u}^{(l-1)}$
  3. UTGIN: $h_{v}^{(l)}=(1+\epsilon)\cdot h_{v}^{(l-1)}+\sum_{u\in\mathcal{N}(v)}h_{u}^{(l-1)}$
@@ 4. UTGraphConv: $h_{v}^{(l)}=\sum_{u\in\tilde{\mathcal{N}}(v)}h_{u}^{(l-1)}$
In general, we consider the case where all nodes have self-loops, i.e. the feature of a node itself is used in the aggregation step. The modified architectures, now simplified, serve as the layers for our untrained experiments. To maintain simplicity and fairness across all methods, we refrain from using edge weights in our models. Further setting $\epsilon=0$ for GINs results in a uniform formula across both models: $h_{v}^{(l)}=\sum_{u\in\tilde{\mathcal{N}}(v)}h_{u}^{(l-1)}$. Henceforth we will refer to both models as UTGIN.
@@
where $\mathbf{H}^{(0)}\in\mathbb{R}^{n\times d}$ is the initial feature matrix, and $\mathbf{H}^{(l)}$ the feature matrix after $l$ iterations of message passing. Following, the definitions of UTMP layers above we have $\mathbf{S}=\tilde{\mathbf{D}}^{-1/2}\tilde{\mathbf{A}}\tilde{\mathbf{D}}^{-1/2}$ for UTGCN and $\mathbf{S}=\tilde{\mathbf{D}}^{-1}\tilde{\mathbf{A}}$ for UTSAGE, where $\tilde{\mathbf{D}}$ is the degree matrix with diagonal entries $\tilde{\mathbf{D}}_{uu}=\sum_{v}\tilde{\mathbf{A}}_{uv}$. Similarly, for UTGIN $\mathbf{S}=\tilde{\mathbf{A}}$. The generalization of UTMP layers to undirected weighted graphs can be obtianed by simply replacing the adjacency matrix and related quantities with their weighted counterparts in the formulation of $\mathbf{S}$.
@@ 3.2 Simplified architectures
Following the construction of Wu et al. for the case of node classification we add a final trained linear layer before the final dot product. We refer to such architectures that include a final trained linear layer after the UTMP layers as ’simplified’ in accordance with Wu et al. (2019) and include an ’S’ in the abbreviations of these models, e.g. SGCN. This results in an architecture where the final node features are given by:
@@
where $\Theta$ is the learned weight matrix of the linear layer. In the case of simplified GNN architectures, the trained linear layer can also be interpreted as a modified positive semi-definite inner product in the form of $<\Theta h_{v}^{l},\Theta h_{u}^{l}>$ where $\Theta$ is the weight matrix of the linear layer.
In the case of link prediction features produced by UTMP layers can actually be used to construct fully untrained architectures that only consist of feature propagation steps followed by an inner product. In practice we found that such architectures based solely on UTMP layers do work well showing that UTMP layers produce highly informative features for link prediction.
@@ 3.3 UTMP layers and path based measures
Building on the formulations of the untrained layers above, in the following we provide a theoretical analysis that relates the inner products of features resulting from untrained message passing layers to pair-wise measures of node similarity that are based on characteristics of paths in the underlying graph. Such path based measures quantify the indirect connection strength between node pairs in the absence of a direct link connecting the nodes. In order to relate path based measures and untrained MP layers we will assume that initial feature vectors are pairwise orthonormal i.e. $<h_{v}^{(0)},h_{u}^{(0)}>=\delta_{u,v}$.
[…]
Now we consider inner products of features after $l$ iterations of message passing which is given by $<h_{u}^{(l)},h_{v}^{(l)}>=(\mathbf{S}^{l}\mathbf{H}^{(0)}\mathbf{H}^{(0)\top}(\mathbf{S}^{l})^{\top})_{uv}$. For orthonormal features the inner products of features reduces to $\mathbf{H}^{(0)}\mathbf{H}^{(0)\top}=\mathbf{I}$ and we obtain the following expression for the inner product of the features after $l$ iterations of UTMP layers:
$<h_{u}^{(l)},h_{v}^{(l)}>=(\mathbf{S}^{l}(\mathbf{S}^{l})^{\top})_{uv}.$
@@
$<h_{u}^{(l)},h_{v}^{(l)}>=\frac{1}{\sqrt{\tilde{d}(u)\tilde{d}(v)}}\sum_{p\in P^{2l}_{uv}}\prod_{i\in[p]}\frac{1}{\tilde{d}_{i}},$
@@
where $[p]$ denotes the path $p$ with the first and last vertices removed. The above expression is equivalent to $\sqrt{P(u\xrightarrow{2l}v)P(v\xrightarrow{2l}u)}$ i.e. the geometric mean of the probabilities that a random walk starting at either $u$ or $v$ to reaches the other in $2l$ steps. Similarly, for UTSAGE we have $\mathbf{S}=\tilde{\mathbf{D}}^{-1}\tilde{\mathbf{A}}$ and the inner product can be expressed in terms of paths in $P^{2l}_{uv}$ as:
$<h_{u}^{(l)},h_{v}^{(l)}>=\sum_{p\in P^{2l}_{uv}}\prod_{i\in p-m(p)}\frac{1}{\tilde{d}_{i}},$
[…]
$<h_{u}^{(l)},h_{v}^{(l)}>=|P^{2l}_{uv}|.$
@@
Although the condition of orthonormality might seem quite restrictive at first glance it applies in many practical settings, though in some cases only approximately. Moreover, orthogonality only needs to be satisfied in the common $l$-neighbourhood of the nodes. One hot encodings are just one example of orthonormal features used in practice. Orthonormality also applies in the case of high dimensional random feature vectors as for sufficiently large dimensions any set of $k$ independent random vectors is almost orthogonal Eaton (2007). Similar results also hold for random high dimensional binary features that are sparsely populated for which the expected value of the inner product of two vectors scales as $O(1/k)$ for dimension $k$.
@@
High dimensional features of empirical data sets also show similar characteristics to their random counterparts. For instance, empirical feature vectors of randomly selected node pairs tend to be approximately orthogonal, notwithstanding the fact that features of connected node pairs can be highly correlated Nt and Maehara (2019), as can be verified experimentally (see Sec.C).
@@
As mentioned before in the case of simplified GNN architectures, the final trained linear layer can be interpreted as a modified positive semi-definite inner product and the orthogonality results for high dimensional random features also apply to such more general inner products. However, note that normality is no longer guaranteed i.e. $<\Theta h_{v},\Theta h_{u}>\sim\delta_{u,v}|h_{u}|^{2}_{\Theta}$.
@@ 3.4 Triadic closure and other path based measures
Triadic closure, also known as transitivity, refers to the tendency for nodes in real-world networks to form connections if they share (many) common neighbors. As such triadic closure has been widely studied as a mechanism that drives link formation in complex real-world networks Rapoport (1953); Holland and Leinhardt (1971). Moreover, node similarity measures that build on triadic closure in social networks have been used for similarity-based link prediction algorithms Lü and Zhou (2011).
Given a pair of nodes $(u,v)$, the tendency of them to be connected due to triadic closure can be quantified by simply counting the number of common neighbours between the two vertices i.e. $T(u,v)=|N(u)\cap N(v)|$ which corresponds to $l=1$ for UTGIN, assuming that $u$ and $v$ are not connected in the graph as is customary in a link prediction scenario. In practice, one might further want to account for the fact that in general nodes with higher degrees also have a larger probability of having common neighbours, for instance by normalizing by the degrees, i.e.: $T_{d}(u,v)=|N(u)\cap N(v)|/\tilde{d}(u)\tilde{d}(v)$, which corresponds to $l=1$ for UTSAGE. One can go one step further and also take into account the degrees of the common neighbours themselves since high degree nodes are by definition common neighbours of more node pairs, for instance by weighing common neighbours according to their degree $T_{n}(u,v)=\frac{1}{\sqrt{\tilde{d}(u)\tilde{d}(v)}}\sum_{i\in N(u)\cap N(v)}\frac{1}{\tilde{d}_{i}}$ which in our case corresponds to UTGCN with $l=1$.
[…]
@@ 4 Experiments and Results
In the following, we provide details on our experimental setup. We evaluate GNN architectures on a variety data sets summarized in Table 3. The data sets cover both attributed graphs where nodes have additional high-dimensional features (Cora small, CiteSeer small, Cora, CoraML, PubMed, CiteSeer, DBLP) and non-attributed graphs that do not contain any node features. Data sources and summary statistics of the data sets can be found in Table 3. We use the area under the Receiver Operator Characteristic curve (ROC-AUC) to link prediction performance for all data sets.
@@ 4.1 Experimental Setup
To ensure a fair comparison among models we maintain the same overall architectures across all experiments where each trainable message passing layer is followed by an Exponential Linear Unit (ELU) and the optimal number of layers for models is determined via hyperparameter search. Upon completion of the message passing layers, we introduce a final linear layer for both trained and simplified models. We also consider untrained (UT) models that do not include this final linear layer and directly take the inner product between the propagated features of the source and target nodes resulting in a parameter-free and hence fully untrained model.
We use Adam Kingma and Ba (2014) as an optimization function and employ binary cross entropy with logits as our loss function. All datasets are transformed by normalizing the node features and randomly splitting the dataset, with $10\%$ allocated to the test set, $5\%$ to the validation set, and the remainder to the training set. Each model configuration is run $10$ times, with the results averaged over these runs. Our training and testing procedures are based on the methodology outlined in pyg (2021), where we perform a new round of negative edge sampling for each training epoch. We limit the maximum number of epochs to 10,000 and also incorporate an early stopping mechanism in our training process by terminating training whenever there is no improvement in the validation set results over a span of $250$ epochs.
For the simplified models we pre-compute node features corresponding to the untrained message passing layers as these do not change during training. We use one-hot encoding as initial node features for the non-attributed datasets.
@@
[…]

### v2

@@ 2. UTSAGE: $h_{v}^{(l)}=\frac{1}{\tilde{d}_{v}}\sum_{u\in\tilde{\mathcal{N}}(v)}h_{u}^{(l-1)}$
  3. UTGIN: $h_{v}^{(l)}=(1+\epsilon)h_{v}^{(l-1)}+\sum_{u\in\mathcal{N}(v)}h_{u}^{(l-1)}$
@@ 4. UTGraphConv: $h_{v}^{(l)}=\sum_{u\in\tilde{\mathcal{N}}(v)}h_{u}^{(l-1)}$
In general, we consider the case where all nodes have self-loops, i.e. the features of the node itself are included in the aggregation step. Further setting $\epsilon=0$ for GINs results in a uniform formula across both models: $h_{v}^{(l)}=\sum_{u\in\tilde{\mathcal{N}}(v)}h_{u}^{(l-1)}$. Henceforth we will refer to both models as UTGIN.
@@
where $\mathbf{H}^{(0)}\in\mathbb{R}^{n\times d}$ is the initial feature matrix, and $\mathbf{H}^{(l)}$ the feature matrix after $l$ iterations of message passing. Following, the definitions of UTMP layers above we have $\mathbf{S}=\tilde{\mathbf{D}}^{-1/2}\tilde{\mathbf{A}}\tilde{\mathbf{D}}^{-1/2}$ for UTGCN and $\mathbf{S}=\tilde{\mathbf{D}}^{-1}\tilde{\mathbf{A}}$ for UTSAGE, where $\tilde{\mathbf{D}}$ is the degree matrix with diagonal entries $\tilde{\mathbf{D}}_{uu}=\sum_{v}\tilde{\mathbf{A}}_{uv}$. Similarly, for UTGIN we have $\mathbf{S}=\tilde{\mathbf{A}}$. The generalization of UTMP layers to undirected weighted graphs can be obtained by simply replacing the adjacency matrix and related quantities with their weighted counterparts in the formulation of $\mathbf{S}$.
@@ 3.2 Simplified architectures
Following the construction of [9] for the case of node classification we add a final trained linear layer before the final dot product. We refer to such architectures that include a final trained linear layer after the UTMP layers as ’simplified’ in accordance with [9] and include an ’S’ in the abbreviations of these models, e.g. SGCN. This results in an architecture where the final node features are given by:
@@
where $\Theta$ is the learned weight matrix of the linear layer. In the case of simplified GNN architectures, the trained linear layer can also be interpreted as a modified positive semi-definite inner product in the form of $\langle\Theta h_{v}^{l},\Theta h_{u}^{l}\rangle$ where $\Theta$ is the weight matrix of the linear layer.
In the case of link prediction features produced by UTMP layers can actually be used to construct fully untrained architectures that only consist of feature propagation steps followed by an inner product. In practice, we found that such architectures based solely on UTMP layers can do surprisingly well in terms of LP performance showing that UTMP layers produce highly informative features.
@@ 3.3 UTMP layers and path based measures
Building on the formulations of the untrained layers above, in the following we provide a theoretical analysis that relates the inner products of features resulting from untrained message passing layers to pair-wise measures of node similarity that are based on characteristics of paths in the underlying graph. Such path based measures offer a way of quantifying the indirect connection strength between node pairs in the absence of a direct link connecting the nodes. In order to relate path based measures and UTMP layers we will assume that initial feature vectors are pairwise orthonormal i.e. $\langle h_{v}^{(0)},h_{u}^{(0)}\rangle=\delta_{u,v}$.
A path of length $l$ is defined as a sequence of $l+1$ vertices $(v_{0},v_{1}\ldots v_{l})$ such that $(v_{i},v_{i+1})\in E$ for all $0\leq i<l$. We denote the space of a set of all paths of length $l$ between nodes $u$ and $v$ as $P^{l}_{uv}$. The number of paths of length $l$ between any $u$ and $v$ is given by the $l^{th}$ power of the adjacency matrix i.e. $|P^{l}_{uv}|=\tilde{\mathbf{A}}^{l}_{uv}$. Note that since we assume self-loops on all vertices $P^{l}_{uv}$ implicitly also includes shorter paths between $u$ and $v$. Similarly, paths of length $l$ between vertices $u$ and $v$ also determine the probability of a random walk starting at $u$ reaching $v$ which is given by $P(u\xrightarrow{l}v)=\sum_{p\in P^{l}_{uv}}\prod_{i\in p-[v]}\frac{1}{\tilde{d}_{i}}$, where p[v]p-[v] denotes that the last vertex ($v$) is not included in the product. The random walk probability can also be expressed in matrix form $P(u\xrightarrow{l}v)=(\tilde{\mathbf{D}}^{-1}\tilde{\mathbf{A}})^{l}_{uv}$.
Now we consider inner products of features after $l$ iterations of message passing which is given by $\langle h_{u}^{(l)},h_{v}^{(l)}\rangle=(\mathbf{S}^{l}\mathbf{H}^{(0)}\mathbf{H}^{(0)\top}(\mathbf{S}^{l})^{\top})_{uv}$. For orthonormal features the inner products of features reduces to $\mathbf{H}^{(0)}\mathbf{H}^{(0)\top}=\mathbf{I}$ and we obtain the following expression for the inner product of the features after $l$ iterations of UTMP layers:
$\langle h_{u}^{(l)},h_{v}^{(l)}\rangle=(\mathbf{S}^{l}(\mathbf{S}^{l})^{\top})_{uv}.$
@@
$\langle h_{u}^{(l)},h_{v}^{(l)}\rangle=\frac{1}{\sqrt{\tilde{d}(u)\tilde{d}(v)}}\sum_{p\in P^{2l}_{uv}}\prod_{i\in[p]}\frac{1}{\tilde{d}_{i}},$
@@
where $[p]$ denotes the path $p$ with the first and last vertices removed. The above expression is equivalent to $\sqrt{P(u\xrightarrow{2l}v)P(v\xrightarrow{2l}u)}$ i.e. the geometric mean of the probabilities that a random walk starting at either $u$ or $v$ reaches the other in $2l$ steps. Similarly, for UTSAGE we have $\mathbf{S}=\tilde{\mathbf{D}}^{-1}\tilde{\mathbf{A}}$ and the inner product can be expressed in terms of paths in $P^{2l}_{uv}$ as:
$\langle h_{u}^{(l)},h_{v}^{(l)}\rangle=\sum_{p\in P^{2l}_{uv}}\prod_{i\in p-m(p)}\frac{1}{\tilde{d}_{i}},$
where $m(p)$ is the midpoint of the path $p$. The above expression is equivalent to $\langle h_{u}^{(l)},h_{v}^{(l)}\rangle=\sum_{i}P(u\xrightarrow{l}i)P(v\xrightarrow{l}i)$ and hence corresponds to the probability that two simultaneous random walks starting at $u$ and $v$, respectively, meet after $l$ steps at some midpoint. Finally, for UTGIN we have $\mathbf{S}=\tilde{\mathbf{A}}$ and hence:
$\langle h_{u}^{(l)},h_{v}^{(l)}\rangle=|P^{2l}_{uv}|.$
@@
Although the condition of orthonormality might seem quite restrictive at first glance it applies in many practical settings, though in some cases only approximately. Moreover, for the above result to hold orthogonality only needs to be satisfied in the common $l$-neighbourhood of the nodes. One example of orthonormal features that are widely used in practice are one hot encodings and orthonormality also applies in the case of high dimensional random feature vectors since for sufficiently large dimensions any set of $k$ independent random vectors is quasi orthogonal [34]. Similar results also hold for random high dimensional binary features that are sparsely populated for which the expected value of the inner product of two vectors scales as $O(1/k)$ for dimension $k$.
@@
High dimensional features of empirical data sets also show similar characteristics to their random counterparts. For instance, empirical feature vectors of randomly selected node pairs tend to be approximately orthogonal, notwithstanding the fact that features of connected node pairs can be highly correlated [35], as can be verified experimentally (see Sec.LABEL:section:_feature_dot_products).
@@
As mentioned before in the case of simplified GNN architectures, the final trained linear layer can be interpreted as a modified positive semi-definite inner product and the orthogonality results for high dimensional random features also apply to such more general inner products. However, note that normality is no longer guaranteed i.e. $\langle\Theta h_{v},\Theta h_{u}\rangle\sim\delta_{u,v}|h_{u}|^{2}_{\Theta}$.
We would like to note that the assumption of orthogonality is a mathematical assumption we use to establish the connection between UTMP layers and path based measures. However, this does not imply that UTMP require orthogonal features to perform well at LP tasks. On the contrary deviations from orthogonality can enhance the LP performance of UTMP for instance when connected nodes tend to have more similar features which holds for many LP benchmarks. We find that the inner products of feature vectors of randomly selected node pairs are in general close to zero. Note that, the feature vectors of all datasets are non-negative as they represent word occurrences. As expected, for connected nodes the inner products of feature vectors tend to be higher reflecting the increased feature similarity.
When initial features deviate from exact orthogonality, we can write their Gram matrix as $\mathbf{H}^{(0)}\mathbf{H}^{(0)\top}=\mathbf{I}+\Delta$, where $\Delta$ captures correlations in the initial features. In this case, the UTMP similarity decomposes as
$\mathbf{S}^{l}\mathbf{H}^{(0)}\mathbf{H}^{(0)\top}(\mathbf{S}^{l})^{\top}=\underbrace{\mathbf{S}^{l}(\mathbf{S}^{l})^{\top}}_{\text{path-based term}}+\underbrace{\mathbf{S}^{l}\Delta(\mathbf{S}^{l})^{\top}}_{\text{feature similarity term}}\,.$
The first term is the pure path-based quantity; the second term reflects the effect of feature similarity transported by $\mathbf{S}^{l}$. Under homophily, this additional term is often beneficial for LP; conversely, large misaligned correlations (large $\Delta$) can introduce noise. Empirically, we find inner products between random node pairs concentrate near zero (Fig. 1), suggesting that $\Delta$ is typically small in our datasets.
@@ 3.4 Triadic closure and other path based measures
Triadic closure, also known as transitivity, refers to the tendency for nodes in real-world networks to form connections if they share (many) common neighbors. As such triadic closure has been widely studied as a mechanism that drives link formation in complex real-world networks [36, 37]. Moreover, node similarity measures that build on triadic closure in social networks have been used for similarity-based link prediction algorithms [38].
Given a pair of nodes $(u,v)$, the tendency of them to be connected due to triadic closure can be quantified by simply counting the number of common neighbours between the two vertices i.e. $T(u,v)=|N(u)\cap N(v)|$ which corresponds to $l=1$ for UTGIN, assuming that $u$ and $v$ are not connected in the graph as is typically the case in an LP setting. In practice, one might further want to account for the fact that in general nodes with higher degrees also have a larger probability of having common neighbours, for instance by normalizing by the degrees, i.e.: $T_{d}(u,v)=|N(u)\cap N(v)|/\tilde{d}(u)\tilde{d}(v)$, which corresponds to $l=1$ for UTSAGE. One can go one step further and also take into account the degrees of the common neighbours themselves since high degree nodes are by definition common neighbours of more node pairs, for instance by weighing common neighbours according to their degree $T_{n}(u,v)=\frac{1}{\sqrt{\tilde{d}(u)\tilde{d}(v)}}\sum_{i\in N(u)\cap N(v)}\frac{1}{\tilde{d}_{i}}$ which in our case corresponds to UTGCN with $l=1$.
Our results also link UTMP layers to other topological similarity measures that are widely used in link prediction heuristics such as the Adamic-Adar (AA) index [39], Resource Allocation (RA) [40], the Katz index [41], rooted PageRank [42] and SimRank [43]. For instance the AA index, given by $AA(u,v)=\sum_{i\in N(u)\cap N(v)}\frac{1}{\log\tilde{d}_{i}}$, and $RA(u,v)=\sum_{i\in N(u)\cap N(v)}\frac{1}{\tilde{d}_{i}}$ differ only slightly from the triadic closure measures we obtained for UTMP layers. Similar results also hold for other path based measures such as rooted PageRank, the Katz index and SimRank which can be defined in terms of power series over paths of different lengths. For instance, SimRank similarity between nodes $u$ and $v$ is defined as $s(x,y)=\sum_{l}P_{uv}(l)\gamma^{l}$ where $P_{uv}(l)$ is the probability that two random walks starting at $u$ and $v$ meet after $l$ steps and $0<\gamma<1$ is a free parameter. Similarly the Katz index is defined as $Katz(u,v)=\sum_{l}\mathbf{A}^{l}_{uv}\gamma^{l}$ and rooted PageRank is defined as $PR(u,v)=(1-\gamma)\sum_{l}\frac{P(u\xrightarrow{l}v)+P(v\xrightarrow{l}u)}{2}\gamma^{l}$ again with $0<\gamma<1$ being a free parameter. Hence, the Katz index is closely related to UTGIN and rooted PageRank is closely related to UTGCN, the main difference being that these measures also include paths of odd length which UTGIN and UTGCN include only indirectly through the inclusion of self loops in their formulation.
@@ 4 Experiments and Results
In the following, we provide details on our experimental setup. We evaluate GNN architectures on a variety data sets that cover both attributed graphs where nodes have additional high-dimensional features (Cora small, CiteSeer small, Cora, CoraML, PubMed, CiteSeer, DBLP) and non-attributed graphs that do not contain any node features. Data sources and summary statistics of the data sets can be found in the Appendix Table 3. We use the area under the Receiver Operator Characteristic curve (ROC-AUC) for the non-attributed datasets and Hits@100 and ROC-AUC for the attributed datasets as our main performance measures. [^1] The code for replicating the results is available at: https://doi.org/10.5281/zenodo.15019863
@@ 4.1 Experimental Setup
To ensure a fair comparison among models we maintain the same overall architectures across all experiments and MP layers. For trainable message passing layers, each layer is followed by an Exponential Linear Unit (ELU) and the optimal number of layers for models is determined via hyperparameter search. Upon completion of the message passing layers, we introduce a final linear layer for both trained and simplified models. We also consider untrained (UT) models that do not include this final linear layer and directly take the inner product between the propagated features of the source and target nodes resulting in a parameter-free and hence fully untrained model. Since the simplified architectures consist of UTMP layers followed by a trainable linear layer, the consideration of UT models which do not include the linear layer also covers all possible ablation studies.
In principle any LP method that uses GNNs as one of its sub-components can also be formulated using UTMP layers. However, in general state of the art methods consist of many sub-components resulting in more complex and computationally demanding experimental setups where the effect of switching from trained to untrained MP layers is difficult to isolate. Therefore, we focus mostly on graph autoencoders in our experiments due to their simplicity, but also consider two versions of NCNC [8]: one which uses trained GNN layers for MP and another that uses UTGNN layers instead, which following our naming convention is denoted as SNCNC.
For the simplified models we precompute node features corresponding to the untrained message passing layers, as these do not change during training. We use one-hot encoding as initial node features for the non-attributed datasets. Further details about the experimental setup can be found in the supplementar material (Sec. C) along with results based on HeaRT [44] evaluation setting (Sec. E) which samples hard negative samples via multiple heuristics.
Table 1: Link Prediction accuracy for attributed networks as measured by Hits@100. Red values correspond to the overall best model for each dataset, and blue values indicate the best-performing model within the same category of message passing layers.
@@
Hits@100 Hits@100 Hits@100 Hits@100 Hits@100 Hits@100 Hits@100
GCN 80.5 ± 1.59 83.0 ± 1.57 79.75 ± 0.74 82.92 ± 1.44 73.64 ± 2.04 81.06 ± 1.3 69.72 ± 1.71
SGCN 84.73 ± 1.46 88.69 ± 0.57 83.93 ± 0.8 87.31 ± 1.23 69.11 ± 1.25 86.02 ± 1.11 64.81 ± 2.09
UTGCN 64.24 ± 3.4 81.07 ± 1.5 37.5 ± 1.5 58.1 ± 1.74 25.35 ± 2.04 69.39 ± 2.22 31.46 ± 1.04
SAGE 75.67 ± 1.29 80.23 ± 1.09 69.07 ± 1.05 78.33 ± 0.85 56.25 ± 0.7 77.14 ± 2.93 64.81 ± 1.66
SSAGE 80.42 ± 1.71 87.22 ± 1.19 74.16 ± 1.47 79.61 ± 1.75 42.14 ± 1.85 83.78 ± 1.61 56.49 ± 2.64
UTSAGE 57.76 ± 1.51 61.85 ± 3.23 30.51 ± 1.57 51.13 ± 1.27 6.6 ± 0.75 69.88 ± 2.15 19.04 ± 2.2
GIN 74.66 ± 1.63 71.16 ± 1.67 69.83 ± 1.07 78.61 ± 1.07 65.3 ± 1.3 74.64 ± 1.61 64.81 ± 1.66
GraphConv 74.7 ± 1.14 74.89 ± 1.59 62.37 ± 1.87 78.66 ± 1.57 62.84 ± 2.1 77.69 ± 1.32 66.59 ± 1.25
SGIN 74.54 ± 1.69 78.71 ± 2.15 73.11 ± 1.03 77.46 ± 1.77 46.21 ± 0.85 78.56 ± 1.31 66.59 ± 1.15
UTGIN 46.73 ± 2.36 61.85 ± 3.23 22.65 ± 1.13 44.79 ± 1.44 22.01 ± 1.71 58.8 ± 6.18 34.29 ± 1.02
NCNC 83.69±3.13 76.37±2.90 84.55±1.14 87.36±1.83 80.72±0.91 87.22±3.61 73.77±0.75
SNCNC 88.72±1.20 93.42±0.78 84.69±1.39 89.81±0.86 81.26±1.59 89.79±1.51 74.23±0.117